A chef is using a mixture of two brands of Italian dressing. The first brand contains 9% vinegar, and the second brand contains 14% vinegar. The chef wants to make 270 milliliters of a dressing that is 12% vinegar. How much of each brand should she use?

Answer by lwsshak3(11628)

A chef is using a mixture of two brands of Italian dressing. The first brand contains 9% vinegar, and the second brand contains 14% vinegar. The chef wants to make 270 milliliters of a dressing that is 12% vinegar. How much of each brand should she use?
**
let x=ml of 9% vineger to be used in mixture
270-x=ml of 14% vineger to be used in mixture
..
9%x+14%(270-x)=12%*270
.09x+37.8-.14x=32.4
.05x=5.4
x=108
270-x=162
ans:
ml of 9% vineger to be used in mixture=108
ml of 14% vineger to be used in mixture=162